# Multicalibration for Censored Survival Data: Towards Universal Adaptability in Predictive Modeling

### v1

@@ 4.1 Property of multicalibration
where $\tau^{(k)}(t),k=1,2$ and $O_{i}$ are defined as in Section 2, and the residual term $R_{N,i}^{(k)}$ is typically negligible. The first and second-order functional derivatives are denoted by $\dot{\phi}^{(k)}(\delta_{O_{i}}-F;t),\ddot{\phi}^{(k)}(\delta_{O_{i}}-F,\delta_{O_{j}}-F;t)$ (see Section LABEL:apxsubsubsec:pseudo-bound in the supplementary material for detailed forms), both of which are linear (bilinear) continuous operators. There is a correspondence between $O_{i}$ and $\delta_{O_{i}}=(Y_{i},N_{i,0},N_{i,1})$, which comprises of three counting processes, where $Y_{i}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\geq s\}$, $N_{i,0}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\leq s,\Delta_{i}=0\}$ and $N_{i,1}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\leq s,\Delta_{i}=1\}$. The limit distribution $F=(H,H_{0},H_{1})$ corresponds to the empirical distribution $F_{N}=N^{-1}\sum_{i=1}^{N}(Y_{i},N_{i,0},N_{i,1})$, with $H(s)=\Pr(\widetilde{T}_{i}\geq s)$, $H_{0}(s)=\Pr(\widetilde{T}_{i}\leq s,\Delta_{i}=0)$, and $H_{1}(s)=\Pr(\widetilde{T}_{i}\leq s,\Delta_{i}=1)$.
@@
The detailed proof of Theorem 3 is deferred to Section LABEL:apxsec:proofs of the supplementary material. The sample size $N$ ought to be large enough so that the empirical average is close enough to the mean. From Theorem 3, we conclude
@@ 5.1 Estimation bias
As expected, due to covariate-shift, Naive estimate of the survival probabilities or restricted mean survival time using the mean estimates from the source population shows a large bias toward the target mean because the unweighted samples are not representative of the true target domain population (Figure LABEL:fig:IPSW-SP-PH-indepLABEL:fig:IPSW-RM-PH-indep). Under mild covariate shift ($q=1$, Figure 2), the ipsw approach alleviates the bias significantly, owing to the sample reweighting that makes nonrandom source samples look like samples drawn from the target domain. ipsw achieves an overall low bias. The proposed post-processed multicalibrated estimators exhibit comparable overall performance.
@@
We observe similar results for the prediction of restricted mean (RM) $\tau^{2}(t)=\mathbb{E}_{\mathcal{D}_{\mathcal{T}}}[T\wedge t]$ in the target domain, as summarized in Figures LABEL:fig:RM-PH-indep-1 and LABEL:fig:RM-PH-indep-3 of Section LABEL:apxsec:simu. We present the relative error because the absolute error grows significantly as $t$ increases. The multicalibrated based approaches learn and calibrate bounded functions that enjoy low prediction errors across all subpopulations, thus outperforming other competitors.
Finally, similar results are observed for estimating both the survival probability and restricted mean survival time in the target domain under moderate shift ($q=2$), as evidenced by Figures LABEL:fig:IPSW-RM-PH-indepLABEL:fig:SP-PH-indep-2 and LABEL:fig:RM-PH-indep-2. In addition to these findings, we also explore other settings, including the accelerated failure hazard model (Tables LABEL:simu:aft-1-LABEL:simu:aft-3) with covariate-dependent Weibull censoring (Tables LABEL:simu:prop-dep-sp-1-LABEL:simu:prop-dep-sp-3) and, more importantly, the scenario with multiple source domains (Tables LABEL:simu:double-sp-1-LABEL:simu:double-rm-3). Our method remains powerful in the sense that it can accommodate multiple source domains, non-proportional hazards, and dependent censoring. Due to the space limit, we refer readers to Section LABEL:apxsec:simu in the supplementary material for more details.
@@ 5.2 The concordance index
We observe that multicalibration also leads to increased C-index as illustrated in Figure 4 and Figure LABEL:fig:Cindex-prop-indep-single-2 under strong covariate-shift, with similar performance under the weak covariate shift. Compared to random forests, the linear models result in higher C-index scores under mild covariate shift, probably attributed to the effective ranking it renders. However, it generally has higher estimation bias as supported by our simulations.

### v2

# Multicalibration for Modeling Censored Survival Data with Universal Adaptability

@@ 4.1 Property of multicalibration
where $\tau^{(k)}(t),k=1,2$ and $O_{i}$ are defined as in Section 2, and the residual term $R_{N,i}^{(k)}$ is typically negligible. The first and second-order functional derivatives are denoted by $\dot{\phi}^{(k)}(\delta_{O_{i}}-F;t),\ddot{\phi}^{(k)}(\delta_{O_{i}}-F,\delta_{O_{j}}-F;t)$ , both of which are linear (bilinear) continuous operators. There is a correspondence between $O_{i}$ and $\delta_{O_{i}}=(Y_{i},N_{i,0},N_{i,1})$, which comprises of three counting processes, where $Y_{i}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\geq s\}$, $N_{i,0}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\leq s,\Delta_{i}=0\}$ and $N_{i,1}(s)={\boldsymbol{1}}\{\widetilde{T}_{i}\leq s,\Delta_{i}=1\}$. The limit distribution $F=(H,H_{0},H_{1})$ corresponds to the empirical distribution $F_{N}=N^{-1}\sum_{i=1}^{N}(Y_{i},N_{i,0},N_{i,1})$, with $H(s)=\Pr(\widetilde{T}_{i}\geq s)$, $H_{0}(s)=\Pr(\widetilde{T}_{i}\leq s,\Delta_{i}=0)$, and $H_{1}(s)=\Pr(\widetilde{T}_{i}\leq s,\Delta_{i}=1)$.
@@
The detailed proof is deferred to the supplementary material. The sample size $N$ ought to be large enough so that the empirical average is close enough to the mean. From Theorem 3, we conclude
@@ 5.1 Estimation bias
As expected, due to covariate-shift, Naive estimate of the survival probabilities or restricted mean survival time using the mean estimates from the source population shows a large bias toward the target mean because the unweighted samples are not representative of the true target domain population (Figure S.8S.9). Under mild covariate shift ($q=1$, Figure 2), the ipsw approach alleviates the bias significantly, owing to the sample reweighting that makes nonrandom source samples look like samples drawn from the target domain. ipsw achieves an overall low bias. The proposed post-processed multicalibrated estimators exhibit comparable overall performance.
@@
We observe similar results for the prediction of restricted mean (RM) $\tau^{2}(t)=\mathbb{E}_{\mathcal{D}_{\mathcal{T}}}[T\wedge t]$ in the target domain, as summarized in Figures S.12 and S.14 of Section S.1. We present the relative error because the absolute error grows significantly as $t$ increases. The multicalibrated based approaches learn and calibrate bounded functions that enjoy low prediction errors across all subpopulations, thus outperforming other competitors.
Finally, similar results are observed for estimating both the survival probability and restricted mean survival time in the target domain under moderate shift ($q=2$), as evidenced by Figures S.9S.11 and S.13. Due to the space limit, we refer readers to Section S.1 in the supplementary material for more details.
@@ 5.2 The concordance index
We observe that multicalibration also leads to increased C-index as illustrated in Figure 4 and Figure S.10 under strong covariate-shift, with similar performance under the weak covariate shift. Compared to random forests, the linear models result in higher C-index scores under mild covariate shift, probably attributed to the effective ranking it renders. However, it generally has higher estimation bias as supported by our simulations.
@@ 6 Application to cardiovascular disease risk prediction
Supplementary Materials for “Multicalibration for Modeling Censored Survival Data with Universal Adaptability”
S.1 Additional Simulations
We have carried out additional simulations that are complementary to the main body of the paper. These simulations were designed to explore different scenarios with varying parameter configurations $(t\in\{5,10,30,50,70\}$, $q=1,2,3)$, thus providing a comprehensive understanding of our model under different conditions. As supplementary to the main content, Figures S.8S.9 depict the improvement of IPSW over naive approach, and Figures S.10-S.11 present results for moderate shift, and Figures S.12 through S.14 display outcome for RM.
Figure S.8: The absolute bias (×$(\times 10^{2})$ for ‘Naive,’ ‘IPSW,’ and ‘IPSW-sub’ in estimating survival probability under varying degrees of covariate shift ($q=1,2,3$), with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results are aggregated from $100$ simulation replications for all subpopulations.
Figure S.9: The relative bias (×$(\times 10^{2})$ for ‘Naive,’ ‘IPSW,’ and ‘IPSW-sub’ in estimating restricted mean under varying degrees of covariate shift ($q=1,2,3$), with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results are aggregated from $100$ simulation replications for all subpopulations.
Figure S.10: C-index for all methods under moderate covariate shift ($q=2$) for various subpopulations. The proportional hazard model with independent censoring following a uniform distribution is considered. Results are based on an average of $100$ simulation replications.
Figure S.11: The absolute bias (×$(\times 10^{2})$ for all methods in estimating survival probability under mild covariate shift ($q=2$) with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results for all subpopulations are presented based on $100$ simulation replications.
Figure S.12: The relative bias (×$(\times 10^{2})$ for all methods in estimating restricted survival mean time under mild covariate shift ($q=1$) with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results for all subpopulations are presented based on $100$ simulation replications.
Figure S.13: The relative bias (×$(\times 10^{2})$ for all methods in estimating the restricted survival mean time under strong covariate shift ($q=2$) with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results for all subpopulations are presented based on $100$ simulation replications.
Figure S.14: The relative bias (×$(\times 10^{2})$ for all methods in estimating the restricted survival mean time under strong covariate shift ($q=3$) with a total sample size of $N=1000$. The survival time adheres to the proportional hazard model, and the censoring is covariate-independent. Results for all subpopulations are presented based on $100$ simulation replications.